Solve the quadratic equation.
4x^2 = 13x + 12
Please help! I broke it down into (2x-12)(2x+1) but using the FOIL method, it turns out to be -20x instead of -13x. Please help! Thanks!

Answer by scott8148(6628)

4x^2-13x-12=0 ... (4x+3)(x-4)=0 ... x=-3/4, x=4